OpenStudy (samigupta8):

if α,β are the roots of equation x^2-2x+3=0 then the equation whose roots are P=α^3-3α^2+5α-2 and Q=β^3-β^2+β+5 is

hartnn (hartnn):

this is long...

hartnn (hartnn):

find sum of roots, alpha + beta and product alpha * beta first

hartnn (hartnn):

then since we have P and Q, we simplify for P+Q and PQ

OpenStudy (samigupta8):

sum is 2 and product is 3

hartnn (hartnn):

yes,

OpenStudy (samigupta8):

if we solve like dat do u think we wud reach to any answer as such

hartnn (hartnn):

P+Q = a^3+b^3 -3a^2 - b^2+5a+b -2+5 no i am not sure, just trying...

OpenStudy (samigupta8):

okk..

hartnn (hartnn):

there is an adjustment involved too

OpenStudy (samigupta8):

which advancement are u talking about

hartnn (hartnn):

P+Q = a^3+b^3 -3a^2 - b^2+5a+b -2+5 =(a^3+b^3) - (a^2+b^2) +(a+b) - ( 2a^2 -4b+6) + 9 given a+b and ab can you find a^3+b^3 a^2+b^2 ?

OpenStudy (samigupta8):

yaa i can find them out

hartnn (hartnn):

then you'll get the value of P+Q similarly try PQ

OpenStudy (samigupta8):

sum =-1

hartnn (hartnn):

ok, for 2a^2 -4b+6 since 'a' is the root a^2-2a+3 = 0 so, 2a^2 = 4a -6 substitute this there...

hartnn (hartnn):

so you'll need the value of a- b too

hartnn (hartnn):

i hope there's some easier method for this...

OpenStudy (samigupta8):

i can find out the value of a-b

hartnn (hartnn):

but did you get what i've done ?

OpenStudy (samigupta8):

yaa

hartnn (hartnn):

i knew we could find a^3 +b^2 , a^2+b^2 thats why i made their groups and written them separately...

OpenStudy (samigupta8):

yaaa

hartnn (hartnn):

oh i made an error

hartnn (hartnn):

P+Q = a^3+b^3 -3a^2 - b^2+5a+b -2+5 =(a^3+b^3) - (a^2+b^2) +(a+b) - ( 2a^2 -4b+6) + 9 5a = a+4a not a+4b so, P+Q = a^3+b^3 -3a^2 - b^2+5a+b -2+5 =(a^3+b^3) - (a^2+b^2) +(a+b) - ( 2a^2 -4a+6) + 9 now 2a62 -4a + 6 =0 you know why ?

hartnn (hartnn):

2a^2 -4a + 6 =0

hartnn (hartnn):

now you will get P+Q = 3 :)

hartnn (hartnn):

try it and see if you get a^3+b^3 = -10 a^2+b^2 = -2

OpenStudy (anonymous):

4PQ=(P+Q)^2-(P-Q)^2 find P-Q and it gives you the value of PQ.

hartnn (hartnn):

nice way to find PQ, directly finding PQ would be nightmare...

OpenStudy (anonymous):

haha..yea! :)

OpenStudy (samigupta8):

do u wanna say that we have to find out p-q also

hartnn (hartnn):

instead of finding PQ, finding P-Q is easier P-Q = a^3-b^3 -3a^2 + b^2+5a-b -2-5 try to group these , like i did ?

hartnn (hartnn):

one group is obviously (a^3-b^3)

OpenStudy (samigupta8):

yaa

OpenStudy (samigupta8):

what about -3a^2+5a+b^2+b

hartnn (hartnn):

-3a^2 = -a^2 -2a^2 5a = a +4a

hartnn (hartnn):

and from original equation we know the value of -2a^2 +4a ...

hartnn (hartnn):

-2a^2 +4a = - (2a^2-4a) ...

OpenStudy (samigupta8):

we have to take value of a-b for that because i m getting eqn as \[a^3-b^3-(2a^2-4a+6)-(a^2+b^2)+a-b+1\]

OpenStudy (samigupta8):

now we have to put that underrot of -8 as value to solve for a^3-b^3 or do u have some other method

hartnn (hartnn):

P-Q = a^3-b^3 -3a^2 + b^2+5a-b -2-5 = (a^3-b^3) + (b^2-a^2) + (a-b) - (2a^2-4a ) -7 no, i don't have another method which will avoid sqrt -8 ...

OpenStudy (samigupta8):

shud we put value of a-b then to get ans

hartnn (hartnn):

try it out, i am not sure..

OpenStudy (samigupta8):

\[(\sqrt{-8})(-2)-(2)(\sqrt{-8})+\sqrt{-8}-1\]

OpenStudy (samigupta8):

it wud give out to be \[(-3\sqrt{-8}-1)\]

OpenStudy (samigupta8):

is it ryt uptil here

hartnn (hartnn):

actually i am getting \((\sqrt{-8})(1)-(2)(\sqrt{-8})+\sqrt{-8}-1\) and the sqrt -8 is getting cancelled, as i suspected! so, P-Q = -1 now use 4PQ=(P+Q)^2-(P-Q)^2 to find PQ and you will get the required equation

hartnn (hartnn):

the \(\large a^3-b^3= (a-b)((a^2+b^2) + ab)\) a^2+b^2 was -2 and ab = 3 so, you'll get only sqrt -8 and not -2sqrt -8 check it out...

OpenStudy (samigupta8):

yaa i got dat